# Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection

### v1

@@ 4. STUDY DESIGN AND RESULTS
In this section, we first describe the study design and setup. Then, we present the motivation, approach, and results of the research questions.
@@
Benchmark Dataset. To answer the RQs, we conducted the experiment on 675 faults across 14 projects from the Defects4J benchmark (V2.0.0) (Just et al., 2014). Defects4J provides a controlled environment to reproduce faults collected from projects of various types and sizes. Defects4J is widely used in prior automated fault localization research (Lou et al., 2021; Sohn and Yoo, 2017; Chen et al., 2022; Zhang et al., 2017). We excluded three projects, JacksonDatabind, JxPath, and Chart, from Defects4J in our study because we encountered many execution errors and were not able to collect test coverage information for them. Table 1 gives detailed information on the projects and faults we use in our study. In total, the faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault). The sizes of the studied projects range from 2K to 90K lines of code. Note that since a fault may have multiple fault-triggering tests, there are more fault-triggering tests than faults.
@@
Implementation and Environment. To collect test coverage data and compute results for baseline techniques, we utilized Gzoltar (Campos et al., 2012), an automated tool that executes tests and gathers coverage information. For the LLM-based components, we employed OpenAI’s GPT-4o mini, a more cost-effective and capable alternative to GPT-3.5 Turbo (OpenAI, 2024). LangChain v0.2 was used to streamline the development of LLM4FL, facilitating the integration of language models with external tools and enhancing the system’s overall functionality (Langchain, 2024a). To implement the self-reflection technique, we leveraged the LangGraph framework, which enabled graph-based reasoning and decision-making processes (Langchain, 2024b). To minimize the variations in the output, we set the temperature parameter to 0 during model inference.
4.1. RQ1: How does LLM4FL perform compared with other fault localization techniques?
@@
Approach. We compare LLM4FL’s fault localization accuracy with five baselines: Ochiai (Abreu et al., 2006), DeepFL (Li et al., 2019), Grace (Lou et al., 2021), DepGraph (Rafi et al., 2024), and AutoFL (Kang et al., 2024).
@@
Ochiai (Abreu et al., 2006) is widely recognized in fault localization research for its high efficiency and accuracy, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024). As such, we use Ochiai to rank the methods during the segmentation process and include it as a baseline for accuracy comparison. DeepFL (Li et al., 2019) is a deep-learning-based fault localization technique that integrates spectrum-based and other metrics such as code complexity, and textual similarity features to locate faults. It utilizes a Multi-layer Perceptron (MLP) model to analyze these varied feature dimensions. We follow the study (Li et al., 2019) to implement DeepFL and include the SBFL scores from 34 techniques, code complexity, and textual similarities as part of the features for the deep learning model. Grace (Lou et al., 2021) is one of the first fault localization techniques based on graph neural networks (GNN) that represents code as a graph and uses a gated graph neural network to rank the faulty methods. Since then, GNN-based techniques have shown superior fault localization accuracy compared to traditional techniques. DepGraph (Rafi et al., 2024) is a GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information in the graph.
AutoFL is a LLM-based fault localization technique. It begins by providing the LLM with information about a failing test and descriptions of specific methods that can be used to navigate the codebase. The LLM then interacts with these methods to gather relevant information, such as covered classes, methods, and code snippets. Scoring and ranking candidate methods (depicted as black rectangles) based on five AutoFL prediction outcomes. The methods are ranked by assigning scores based on multiple runs of AutoFL. In each run, a method’s score is the inverse of the total number of predicted methods, and these scores are averaged across all runs. Methods are then ranked in descending order of their average scores, with earlier predictions used to break any ties. While the original paper used OpenAI’s GPT-3.5-turbo-0613 model for their experiments, for our evaluation, we are using the latest lightweight GPT-4o mini model to perform AutoFL’s experiments.
Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project. The result continues in Table 3.
        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
        Cli (39)         Ochiai         3.0         5.0         10.0         18.0
        DeepFL         11.0         21.0         24.0         28.0
        Grace         14.0         24.0         26.0         30.0
        DepGraph         17.0         24.0         27.0         34.0
        AutoFL         12.0         19.0         19.0         20.0
        LLM4FL         16.0         21.0         23.0         24.0
        Closure (174)         Ochiai         20.0         39.0         70.0         72.0
        DeepFL         46.0         61.0         92.0         99.0
        Grace         51.0         78.0         102.0         121.0
        DepGraph         60.0         99.0         126.0         148.0
        AutoFL         45.0         58.0         65.0         82.0
        LLM4FL         52.0         77.0         102.0         118.0
        Codec (18)         Ochiai         3.0         12.0         17.0         17.0
        DeepFL         5.0         10.0         12.0         16.0
        Grace         6.0         11.0         13.0         17.0
        DepGraph         7.0         10.0         14.0         16.0
        AutoFL         12.0         14.0         14.0         16.0
        LLM4FL         9.0         13.0         13.0         13.0
        Collections (4)         Ochiai         1.0         1.0         2.0         2.0
        DeepFL         1.0         1.0         2.0         2.0
        Grace         1.0         1.0         2.0         2.0
        DepGraph         1.0         2.0         2.0         2.0
        AutoFL         1.0         1.0         1.0         1.0
        LLM4FL         1.0         1.0         1.0         1.0
        Compress (47)         Ochiai         5.0         12.0         17.0         29.0
        DeepFL         22.0         27.0         31.0         38.0
        Grace         23.0         29.0         34.0         42.0
        DepGraph         25.0         33.0         36.0         45.0
        AutoFL         23.0         33.0         34.0         35.0
        LLM4FL         23.0         32.0         34.0         34.0
        Csv (16)         Ochiai         3.0         8.0         10.0         12.0
        DeepFL         7.0         8.0         9.0         11.0
        Grace         6.0         8.0         10.0         12.0
        DepGraph         8.0         9.0         12.0         13.0
        AutoFL         5.0         11.0         12.0         14.0
        LLM4FL         8.0         10.0         10.0         10.0
        Gson (16)         Ochiai         4.0         9.0         9.0         12.0
        DeepFL         8.0         11.0         12.0         12.0
        Grace         11.0         13.0         14.0         15.0
        DepGraph         14.0         15.0         16.0         16.0
        AutoFL         5.0         7.0         10.0         10.0
        LLM4FL         11.0         14.0         14.0         14.0
@@
Table 3. Continued from Table 2. Fault localization accuracy in terms of Top-1, 3, 5, and 10. The numbers in the parenthesis show the number of faults in each project.
        Project (# faults)         Techniques         Top-1         Top-3         Top-5         Top-10
        JacksonCore (26)         Ochiai         6.0         11.0         13.0         14.0
        DeepFL         5.0         5.0         9.0         10.0
        Grace         9.0         13.0         14.0         15.0
        DepGraph         12.0         15.0         15.0         16.0
        AutoFL         10.0         17.0         17.0         17.0
        LLM4FL         12.0         13.0         15.0         15.0
        JacksonXml (6)         Ochiai         0.0         0.0         0.0         0.0
        DeepFL         3.0         3.0         4.0         5.0
        Grace         3.0         3.0         4.0         5.0
        DepGraph         4.0         5.0         5.0         5.0
        AutoFL         2.0         2.0         2.0         3.0
        LLM4FL         4.0         4.0         4.0         4.0
        Jsoup (93)         Ochiai         15.0         40.0         48.0         57.0
        DeepFL         33.0         39.0         46.0         49.0
        Grace         40.0         64.0         72.0         77.0
        DepGraph         53.0         73.0         78.0         83.0
        AutoFL         36.0         52.0         52.0         54.0
        LLM4FL         41.0         56.0         60.0         60.0
        Lang (64)         Ochiai         25.0         45.0         51.0         59.0
        DeepFL         42.0         53.0         55.0         57.0
        Grace         43.0         53.0         57.0         58.0
        DepGraph         48.0         55.0         60.0         61.0
        AutoFL         40.0         57.0         60.0         60.0
        LLM4FL         48.0         55.0         58.0         58.0
        Math (106)         Ochiai         23.0         52.0         62.0         82.0
        DeepFL         52.0         81.0         90.0         95.0
        Grace         64.0         79.0         92.0         97.0
        DepGraph         72.0         92.0         97.0         102.0
        AutoFL         53.0         81.0         87.0         94.0
        LLM4FL         68.0         87.0         92.0         94.0
        Time (26)         Ochiai         6.0         12.0         13.0         16.0
        DeepFL         12.0         15.0         18.0         20.0
        Grace         11.0         16.0         20.0         21.0
        DepGraph         17.0         20.0         21.0         22.0
        AutoFL         13.0         18.0         21.0         22.0
        LLM4FL         14.0         21.0         22.0         23.0
        Mockito (38)         Ochiai         7.0         14.0         18.0         23.0
        DeepFL         10.0         18.0         23.0         26.0
        Grace         16.0         24.0         26.0         29.0
        DepGraph         21.0         29.0         32.0         34.0
        AutoFL         18.0         23.0         29.0         29.0
        LLM4FL         21.0         22.0         26.0         27.0
        Total (675)         Ochiai         121.0         260.0         340.0         413.0
        DeepFL         257.0         353.0         427.0         468.0
        Grace         298.0         416.0         486.0         541.0
        DepGraph         359.0         481.0         541.0         597.0
        AutoFL         275.0         393.0         423.0         457.0
        LLM4FL         328.0         426.0         474.0         495.0
@@
Results. LLM4FL outperforms the LLM-based baseline, AutoFL, by achieving a much higher Pass@1 (328 v.s. 275). Table 2 and 3 show the fault localization results of LLM4FL and the baseline techniques. Between the two LLM-based techniques, LLM4FL achieves a better Top@N than AutoFL when N is 1, 3, 5, and 10. In the Top-1 metric, LLM4FL locates the correct fault in 328 cases, compared to AutoFL’s 275, representing a 19.27% improvement. Similarly, in the Top-3 and Top-5 metrics, LLM4FL achieves scores of 426 and 474, respectively, outperforming AutoFL’s results of 393 and 423. Even in the broader Top-10 metric, LLM4FL shows an 8.32% advantage, with a score of 495 compared to AutoFL’s 457. These numbers highlight LLM4FL’s enhanced ability to pinpoint faulty methods more accurately and efficiently, reinforcing its superiority over AutoFL for fault localization tasks.
LLM4FL shows higher Top-1 and Top-3 compared to all other non-LLM-based techniques, except for DepGraph. For the Top-1 metric, LLM4FL scores 328, which is 171.07% higher than Ochai’s score of 121, 27.63% higher than DeepFL’s score of 257, and 10.07% better than Grace’s result of 298. One exception is DepGraph, which achieves a Top-1 of 359, 8.64% higher than LLM4FL. As the range expands to Top-3 and beyond, LLM4FL continues to demonstrate its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace. It is important to note that DeepFL, Grace, and DepGraph are supervised methods trained and evaluated using leave-one-out cross-validation, following the original study protocols. Despite LLM4FL being a zero-shot approach without task-specific training, its ability to perform competitively with these supervised techniques further highlights its effectiveness. The fact that LLM4FL achieves strong results without the need for extensive training data or cross-validation underscores its potential and sheds light on future research directions, especially when training models on specific data sets may not always be feasible.
LLM4FL outperforms the other LLM-based technique, AutoFL, by achieving a 19.27% higher Top-1 and delivering competitive results compared to supervised techniques like DeepFL and Grace, even without task-specific training. The findings highlight the potential of using LLM4FL for zero-shot fault localization.
4.2. RQ2: How do different components affect LLM4FL’s fault localization accuracy?
Motivation. In this research question, we study the influence of each individual component within LLM4FL on its fault localization performance. LLM4FL employs a combination of advanced techniques, including coverage splitting to handle token limitations, prompt chaining for fault navigation, self-reflection for iterative refinement of the agents’ decisions, and chain-of-thought reasoning for re-ranking suspicious methods. Each of these components plays a distinct role in the overall process. To study their effects, we investigate how the removal of any one of these components affects the LLM4FL’s ability to accurately localize faults. By isolating the effects of each technique, we can better understand which components are critical to maintaining or improving the overall accuracy of the LLM4FL and provide insights to improving future LLM-based fault localization techniques.
Approach. To evaluate the impact of each component, we designed four experimental configurations: LLM4FLw/o CodeNavigation, LLM4FLw/o Divide&Conquer, LLM4FLw/o Reflection, and LLM4FLw/o CoT. For each configuration, one component is systematically removed, and we compare the result with the full-featured baseline of LLM4FL. This allows us to directly measure how much each component contributes to the fault localization process.
@@ 4.2. RQ2: How do different components affect LLM4FL’s fault localization accuracy?
LLM4FLw/o CodeNavigation removes the prompt chaining mechanism, which essentially means the agents no longer collaborate do fault navigation through multiple rounds of reasoning. Instead, the LLM4FL uses a single prompt to perform fault localization without iterative agent communication and fault navigation. This configuration tests whether the multi-step agent collaboration improves the ranking and selection of faulty methods or if a single round of communication is sufficient.
@@
LLM4FLw/o Divide&Conquer removes the grouping of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups. Coverage segmentation addresses token limitations in LLMs, so removing it explores the impact of feeding the full dataset to the agents in one step. We aim to see how handling large amounts of data in a single input influences the fault localization result, as it may overwhelm the model or reduce precision.
@@
LLM4FLw/o Reflection removes the self-reflection technique, which is used to allow agents to review and refine their initial candidate selection and ranking. Without this self-reflection step, the agents rely solely on their initial assessments without iterative improvements.
LLM4FLw/o CoT disables the chain-of-thought reasoning process, which asks the agents to think deeper by asking LLMs to generate potential fixes and re-rank the suspicious methods. The chain-of-thought approach is designed to enhance logical reasoning and ensure that intermediate results are critically evaluated and refined.

### v2

@@ 4. STUDY DESIGN AND RESULTS
In this section, we first describe the study design and setup. Then, we present the answers to the research questions.
@@
Benchmark Dataset. To answer the RQs, we conducted the experiment on 675 faults across 14 projects from the Defects4J benchmark (V2.0.0) (Just et al., 2014). Defects4J provides a controlled environment to reproduce faults collected from various types and sizes of projects. Defects4J is widely used in prior automated fault localization research (Lou et al., 2021; Sohn and Yoo, 2017; Chen et al., 2022; Zhang et al., 2017). We excluded three projects, JacksonDatabind, JxPath, and Chart, from Defects4J in our study because we encountered many execution errors and were not able to collect test coverage information for them. Table 2 gives detailed information on the projects and faults we use in our study. The faults have over 1.4K fault-triggering tests (i.e., failing tests that cover the fault). The sizes of the studied projects range from 2K to 90K lines of code. Note that since a fault may have multiple fault-triggering tests, there are more fault-triggering tests than faults.
@@
Implementation and Environment. To collect test coverage data and compute results for baseline techniques, we utilized Gzoltar (Campos et al., 2012), an automated tool that executes tests and gathers coverage information. For the LLM-based components, we employed OpenAI’s gpt-4o-mini-2024-07-18, a more cost-effective yet capable LLM (OpenAI, 2024). We used LangChain to develop LLM4FL (Langchain, 2024). We designed the prompts to be concise to minimize token usage and to allow more room for analysis-related information and code. To reduce the variations in the output, we set the temperature parameter to 0 during model inference.
RQ1: How does LLM4FL perform compared with other fault localization techniques?
@@
Approach. We compare LLM4FL’s fault localization accuracy with five baselines representing different methodological families: Ochiai (statistical) (Abreu et al., 2006), DeepFL (deep neural network) (Li et al., 2019), Grace (graph neural network) (Lou et al., 2021), DepGraph (graph neural network) (Rafi et al., 2024), AutoFL (LLM) (Kang et al., 2024), and AgentFL/SoapFL (Qin et al., 2024) (LLM).
Ochiai (Abreu et al., 2006) is a widely recognized statistical fault localization technique known for its high efficiency, making it a common baseline for comparison (Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021; Rafi et al., 2024). As such, we use Ochiai to rank the methods during the segmentation process and include it as a baseline for accuracy comparison.
@@
DeepFL (Li et al., 2019) is a deep-learning-based fault localization technique that integrates spectrum-based and other metrics such as code complexity, and textual similarity features to locate faults. It utilizes a Multi-layer Perceptron (MLP) model to analyze these varied feature dimensions. We follow the study (Li et al., 2019) to implement DeepFL and include the SBFL scores from 34 techniques, code complexity, and textual similarities as part of the features for the deep learning model. Grace (Lou et al., 2021) utilized graph neural networks (GNN). It represents code as a graph and uses a gated graph neural network to rank the faulty methods. DepGraph (Rafi et al., 2024) is another GNN-based technique that further improves Grace by enhancing code representation in a graph using interprocedural call graph analysis for graph pruning and integrating historical code change information.
AutoFL (Kang et al., 2024) is an LLM-based fault localization approach that provides the LLM with a failing test and method descriptions to gather relevant coverage data, then repeats the process to assign inverse scores and rank candidate methods by averaging results across runs. SoapFL (also known as AgentFL(Qin et al., 2024) organizes fault localization into structured phases, including test failure comprehension, code navigation, and iterative LLM-based scoring, leveraging enhanced method documentation to refine suspicious methods. While both AutoFL and SoapFL used OpenAI’s GPT-3.5 for their experiments, for our evaluation, we adapted both techniques to use the same LLM version as LLM4FL (i.e., gpt-4o-mini-2024-07-18) for comparison.
@@
Table 3. Fault localization accuracy (Top-1, 3, 5, and 10) for 675 faults from Defect4J V2.0.0. The numbers in the parenthesis show the percentage difference compared to LLM4FL.
Techniques Top-1 Top-3 Top-5 Top-10
Ochiai 121 (169.42%) 260 (63.46%) 340 (39.41%) 413 (20.10%)
DeepFL 257 (26.85%) 353 (20.40%) 427 (11.01%) 468 (5.98%)
Grace 298 (9.40%) 416 (2.16%) 486 (-2.47%) 541 (-8.32%)
DepGraph 359 (-9.19%) 481 (-11.64%) 541 (-12.38%) 597 (-16.92%)
AutoFL 275 (18.55%) 393 (8.14%) 423 (12.06%) 457 (8.53%)
SoapFL/AgentFL 311 (4.82%) 414 (2.66%) 455 (4.18%) 478 (3.77%)
LLM4FL 326 425 474 496
@@
Results. LLM4FL outperforms the LLM-based baselines, AutoFL and AgentFL, by achieving a 185.55% and 4.82% improvement in Pass@1, respectively. Tables 3 show the fault localization results of LLM4FL and the baseline techniques. Among the three LLM-based techniques, LLM4FL achieves a better Top@N across all values of N. In the Top-1 metric, LLM4FL locates the correct fault in 326 cases, compared to AutoFL’s 275 and AgentFL’s 311, representing an 18.55% and 4.82% improvement, respectively. Similarly, in Top-3, Top-5, and Top-10, LLM4FL an improvement between 8.14% to 12.06% over AutoFL and 2.66% to 4.18% for AgentFL. These numbers highlight LLM4FL’s ability to pinpoint faulty methods more accurately. For cost, we compute the total dollars spent by multiplying each approach’s token usage by the per-million-token price (e.g., \$0.150 for input, \$0.600 for output). SOAPFL averages about \$0.055 per bug, AutoFL averages about \$0.065, while LLM4FL is at around \$0.050 per bug making our approach comparably cost-effective.
LLM4FL shows higher Top-1 and Top-3 compared to most other non-LLM-based techniques. For the Top-1 metric, LLM4FL scores 326, which is 169.42% higher than Ochia’s score of 121, 26.84% higher than DeepFL’s score of 257, and 9.39% better than Grace’s result of 298. One exception is DepGraph, which achieves a Top-1 of 359, 8.64% higher than LLM4FL. As the range expands to Top-3 and beyond, LLM4FL demonstrates its robustness, significantly outperforming DeepFL and maintaining strong performance alongside Grace. LLM-based techniques have several advantages over traditional techniques such as DepGraph. First, techniques like LLM4FL leverage pre-trained LLMs in a zero-shot setting, which can be easily applied to systems with insufficient training data. Second, techniques like DepGraph require additional model training that can take days. LLM-based techniques leverage generic pre-trained LLMs without specific fine-tuning or re-training. Finally, LLM-based techniques can explain the decision, which helps with adaption (Kang et al., 2024; Qin et al., 2024). Hence, LLM-based techniques have a strong potential to enhance fault localization by offering greater adaptability, reducing training overhead, and providing interpretable explanations.
LLM4FL achieves a higher Top-1 compared to other LLM-based techniques, AutoFL and AgentFL, by 18.55% and 4.82%, respectively. It also has competitive results compared to supervised techniques like DeepFL and Grace by leveraging a pre-trained LLM in a zero-shot setting.
RQ2: How do different components in LLM4FL affect the fault localization accuracy?
Motivation. LLM4FL employs a number of components, and each of these components plays a distinct role in the overall process. Hence, in this RQ, we conduct an ablation study by removing each component separately and study their impact on fault localization accuracy. The findings may inspire future studies on adapting the components for similar tasks.
Approach. To evaluate the impact of each component, we designed four different configurations:
@@
LLM4FLw/o CodeNav removes the code navigation mechanism, which essentially means the Debugger agent no longer does fault navigation by retrieving the call graphs. Instead, the LLM4FL uses a single prompt to perform fault localization without fault navigation. This configuration tests whether using the caller-callee information improves the ranking and selection of faulty methods or if a single round of analysis is sufficient.
@@
LLM4FLw/o Division removes the order-aware division of the covered methods, giving the agents the entire coverage data at once instead of dividing it into smaller, manageable groups. Coverage segmentation addresses token limitations in LLMs, so removing it explores the impact of feeding the full dataset to the agents in one step. We aim to see how handling large amounts of data in a single input influences the fault localization result, as it may overwhelm the model or reduce precision.
@@
LLM4FLw/o Reflexion removes the verbal reinforcement learning technique, which is used to allow agents to review and refine their initial ranking. Without this step, the agents rely solely on their initial assessments without iterative improvements.